OpenStudy (anonymous):

please help :( Use the function y=-2(x+2)(2x+8). What are the x- intercepts of the functions graph. What are the coordinates of the graphs vertex

OpenStudy (anonymous):

we want to get this into vertex form first. vertext form looks like: y = a(x-h)^2 + k so, we FOIL to get rid of the ( ) ( ) y=-2(x+2)(2x+8). y=-2(2x^2+8x+4x+16) y=-2(2x^2+12x+16) now we pull out the constant of 16 from the ( ) y=-2(2x^2+12x) +16(-2) y=-2(2x^2+12x) +(-32) now we complete the square for the (2x^2 + 12x) portion of the equation, can you do this for me?